# Classification of unimodal isolated complete intersection singularities in positive characteristic

### v2

@@ 6.2. The classification of unimodal
If $f$ is of the form 6.4 with $v=4,s>4$, then $f\sim N_{\lambda},N^{2}_{\lambda}$ or $P_{r,\infty},P_{\infty,s},P_{r,s,\lambda}$.
@@
If $r\geq 4,$ then let $g=(x^{3},x^{2}y+y^{4})$ be the weighted 0-jet of $f$ w.r.t. $(a,d)$, where $a=(3,2),d=(9,8)$. Let $T=span\langle(0,xy^{3})\rangle$, then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$, which means $f\sim(x^{3},x^{2}y+y^{4}+\lambda xy^{3})$ for $\lambda\in F$.
If $r=3$, then let $g=(x^{3}+a(0)xy^{3},x^{2}y+y^{4}),\ a(0)\in F^{\times}$ be the weighted 0-jet of $f$ w.r.t. $(a,d)$, where $a=(3,2),d=(9,8)$. If $a(0)\neq 1$, then $F_{a,d}^{1}R^{2}\subset\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$ automatically holds, hence $f\sim(x^{3}+\lambda xy^{3},x^{2}y+y^{4}),\ \lambda\in F^{\times}$.
@@
If $r\geq 4,$ then let $g=(x^{3}+by^{5},x^{2}y)$ with $b\in F^{\times}$ be the weighted 0-jet of $f$ w.r.t. $(a,d)$, where $a=(5,3),d=(15,13)$. Same to Proposition 5.14, $f\sim(x^{3}+y^{5},x^{2}y+\lambda y^{5}),\ \lambda\in\{0,1\}$.
If $r=3$, let $g=(x^{3}+axy^{3},x^{2}y)$ with $a\in F^{\times}$ be the weighted 0-jet of $f$ w.r.t. $(a,d)$, where $a=(3,2),d=(9,8)$. Same to Proposition 5.16, we have $f\sim(x^{3}+xy^{3},x^{2}y+y^{t}),\ t\geq 5$. ∎
@@
In this case we have $r=3$ by Proposition 6.8. Under the assumption $s,v>4$, we can choose $g=(x^{3}+axy^{3},x^{2}y)$ be the weighted 0-jet of $f$ w.r.t. $(a,d)$, where $a=(3,2),d=(9,8)$. Let $T=span\langle(0,y^{l})|\ l\geq 5\rangle$, then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$. Therefore $f\sim(x^{3}+xy^{3},x^{2}y+e(y)y^{t})$. Using $r,s$-trick same as Proposition 5.16, we have $f\sim(x^{3}+xy^{3},x^{2}y+y^{t}),\ t\geq 5$. ∎
@@ 7. The classification of order 3 when $charF=3$
7.1. The classification of 3-jet
Same as in Section 5, for $f=(f_{1},f_{2})$ with $\mathrm{ord}(f_{1})=3$, we have $j_{3}(f_{1})\sim(x^{3},0),(x^{2}y,0)$ or $x^{3}+xy^{2}$.
@@ 7.1. The classification of 3-jet
II.4 If $a,b,c\neq 0,$ change the notations, $j_{3}(f)\sim(x^{2}y,y^{3}+ax^{3}+bxy^{2}).$ Applying the automorphism $\phi(x)=x,\phi(y)=y+\alpha x$, where $\alpha$ is a nonzero root of $\alpha^{3}-b\alpha^{2}+a=0$. In this case $\alpha\neq b$ since $a\neq 0$. Then $j_{3}(f)\sim(x^{2}y+\alpha x^{3},y^{3}+(\alpha^{3}+b\alpha^{2}+a)x^{3}+bxy^{2}+2b\alpha x^{2}y)\sim(x^{2}y+\alpha x^{3},y^{3}+bxy^{2}+2b\alpha(x^{2}y+\alpha x^{3}))\sim(x^{2}y+\alpha x^{3},y^{3}+bxy^{2})\sim(x^{3}+x^{2}y,y^{3}+\lambda xy^{2}),\lambda\neq 1$ since $\alpha\neq b$.
III. $j_{3}(f_{1})\sim x^{3}+xy^{2}.$ We have $j_{3}(f)\sim(x^{3}+xy^{2},axy^{2}+bx^{2}y+cy^{3})$. All the discussion is same with in Section 5, except $3=0$ in 5.5, with nothing changes.
@@
(7.1) $\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},y^{3}+xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{2}y,y^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda xy^{2})(\lambda\neq 1),\\ &(x^{3}+xy^{2},0).\end{split}$
@@ 7.2. The classification of unimodal
$u=u_{0}+u_{1}y+\dots,\ v=v0+v_{1}y+\dots,$ $3\mid 2r-2s$
@@
Since most of the discussion are the same as before, we make a table to present the results. See Table LABEL:table_char3-process.
@@
When $j_{3}(f)\sim(x^{3},xy^{2}),(x^{2}y,y^{3}+xy^{2}),(x^{3},x^{2}y)$, we need further discuss.
@@
If $3\mid r$, reset notations, write $f\sim(e(y)x^{3}+y^{r},xy^{2})\sim(x^{3}+e_{1}x^{3}y+y^{r},xy^{2})$, where $e_{i}\in F$. Then $f\sim\widetilde{M}^{3}_{r,e_{1}}$.
@@
If $3\mid r$, similar as above, $f\sim(y^{3}+xy^{2}+\lambda xy^{3}+x^{r},x^{2}y)\sim\widetilde{K}^{3}_{r,\lambda}$ with $\lambda\in F$.
@@
Here we omit the discussion same as Proposition 5.12 to 5.16 and give the result directly in Table LABEL:eqtable.
@@
$s=4$ $(x^{3}+y^{4},x^{2}y)$ (4,3;(4,3;( 4 , 3 ; $(0,y^{4})$ $(x^{3}+y^{4},x^{2}y+\mu y^{4}),$
12,11)12,11)12 , 11 ) $\mu\in\{0,1\}$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ 00 $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$
$r=3$ $\lambda\notin\{0,1\}$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ $(0,y^{t}),t\geq 5$ $(x^{3}+xy^{3},x^{2}y+y^{t})$
$r=3$ $\lambda=0$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ $(0,xy^{r}),(0,y^{s})$ $(x^{3}+xy^{3}+uxy^{r}+vy^{s},$
$r=3$ $\lambda=1$ $x^{2}y+y^{4})$
@@ 8. Check the modality
Let $T^{1,sec}(f)=R^{m}\bigg{/}\bigg{(}\langle f_{1},\dots,f_{m}\rangle\cdot R^{m}+\mathfrak{m}\langle\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}\rangle\bigg{)}$. Choose $g_{1},\dots,g_{d}$ to be a $F$-basis of $T^{1,sec}(f)$. T.H. Pham, Gerhard Pfister, and G.-M. Greuel has shown in [PPG24] that
$F_{t}(x)=F(x,t_{1},\dots,t_{d})=f+t_{1}g_{1}+\dots+t_{d}g_{d}$
represents a formally semiuniversal deformation of $f$. If $F(x,t)$ is equivalent to a family of ICIS of at most one parameter for $t\in F^{d}$, then $f$ is unimodal. Here we check $f\sim l_{q,\lambda}\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q})$ in Table LABEL:table-ord2 as an example:
@@
Note that $T^{1,sec}=(\mathfrak{m}R^{2})/\widetilde{T}_{f}(\mathcal{K}f)$. In the proof of (3.b.ii.2.3.3.1), we have shown that $(0,xy^{q+1})\in\widetilde{T}_{f}(\mathcal{K}f)$. That means $(0,y^{q+3})\in\widetilde{T}_{f}(\mathcal{K}f)$. We also have $(0,y^{q+2})$ generates $(0,xy^{q})$, $(y^{3},0)$ and $(0,y^{q+1})$ generate $(0,xy^{q-1})$. Then we add $(0,x),(0,xy),\dots,(0,xy^{q-2})$ as generators. These generators form a basis of $T^{1,sec}(f)$.
@@
We consider $G=(G_{1},G_{2})=F(x,t)=f+t_{1}g_{1}+\dots+t_{2q+1}g_{2q+1}$.
@@
(8.1) $G$ $\sim(x^{2}+y^{4},y^{q+2}+\lambda y^{q}+t_{5}y^{4}+\dots+t_{q+3}y^{q+2}+t_{q+5}xy^{2}+\dots+t_{2q+1}xy^{q-2})$
$\sim(x^{2}+y^{4},\sum_{i\geq u}t_{i+1}y^{i}+x\sum_{j\geq v}t_{j+q+3}y^{j}).$
Comparing with 3.1 and Proposition 3.6, it corresponds to the case $\alpha=1,\ s=4,\ u=t\geq 4,\ v=q\geq 2$, where $\alpha,s,t,q$ is in the sense of 3.1.
By Proposition 3.6 (3.a.1), in most of cases, $G$ is at most unimodal. The only unsure case is that there exists $q^{\prime}<t<t^{\prime}$ such that
@@
$\widetilde{l}_{q,t,t^{\prime}}$ $(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+uxy^{t}+xy^{t^{\prime}})$ $\lambda^{2}=-1,q\geq 3,$ $q\leq 2p+3$
$t\geq q+1,t^{\prime}\geq t+1,$
@@
for general $a,b,c,d\in F$, then we can ensure that the singularities given in Table LABEL:table-ord2-mod1 are the only unimodal ICIS. Conversely, if all singularities of the form $\widetilde{l}_{q,t,t^{\prime}}$ are equivalent (or at least can be presented as an one parameter family), then all the singularities given in Table LABEL:table-ord2 are unimodal. Unfortunately, we cannot judge this equivalence yet. We post it as a conjecture.
@@
That means, every singularities in Table LABEL:table-ord3 is unimodal. This is because the bimodular one must have a deformation to $\widetilde{P}_{r,s}$. But $(xy^{k},0),(y^{l},0)\ (k\geq 4,l\geq 5)$ do not exist in a basis of $T^{1,sec}(f)$ at the same time for every $f$. Then such a deformation do not exist.
For $ord(f)=3,\ charF=2$ case, see Table LABEL:table-ord3-char2-mod1:
@@
$\widetilde{M}_{r}^{2}$ $(x^{3}+y^{r}+ey^{l},x^{2}y+xy^{2}),$ $r\geq 4,\ r\ is\ odd$ $l\leq 7$
where $e=e_{0}+e_{1}y^{2}+e_{2}y^{4}+\dots$ $l>r,\ l\ is\ even$
@@
For $ord(f)=3,\ char(F)=3$ case, we can check that every singularities in Table LABEL:table-ord3-char3 is unimodal.

### v3

@@ II $(x^{3},y^{3}+x^{2}y)$
[…]
@@
If $r\geq 4,$ then let $g=(x^{3},x^{2}y+y^{4})$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where $a=(3,2),d=(9,8)$. Let $T=span\langle(0,xy^{3})\rangle$, then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$, which means $f\sim(x^{3},x^{2}y+y^{4}+\lambda xy^{3})$ for $\lambda\in F$.
If $r=3$, then let $g=(x^{3}+a(0)xy^{3},x^{2}y+y^{4}),\ a(0)\in F^{\times}$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where $a=(3,2),d=(9,8)$. If $a(0)\neq 1$, then $F_{a,d}^{1}R^{2}\subset\widetilde{T}_{g}(F^{1}\mathcal{K}g)$ automatically holds, hence $f\sim(x^{3}+\lambda xy^{3},x^{2}y+y^{4}),\ \lambda\in F^{\times}$.
@@
If $r\geq 4,$ then let $g=(x^{3}+by^{5},x^{2}y)$ with $b\in F^{\times}$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where $a=(5,3),d=(15,13)$. Same to Proposition 5.14, $f\sim(x^{3}+y^{5},x^{2}y+\lambda y^{5}),\ \lambda\in\{0,1\}$.
If $r=3$, let $g=(x^{3}+axy^{3},x^{2}y)$ with $a\in F^{\times}$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where $a=(3,2),d=(9,8)$. Same to Proposition 5.16, we have $f\sim(x^{3}+xy^{3},x^{2}y+y^{t}),\ t\geq 5$. ∎
@@
In this case we have $r=3$ by Proposition 6.8. Under the assumption $s,v>4$, we can choose $g=(x^{3}+axy^{3},x^{2}y)$ be the weighted 0-jet of $f$ with respect to $(a,d)$, where $a=(3,2),d=(9,8)$. Let $T=span\langle(0,y^{l})|\ l\geq 5\rangle$, then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$. Therefore $f\sim(x^{3}+xy^{3},x^{2}y+e(y)y^{t})$. Using $r,s$-trick same as Proposition 5.16, we have $f\sim(x^{3}+xy^{3},x^{2}y+y^{t}),\ t\geq 5$. ∎
@@ 7. The classification of order 3 when $charF=3$
7.1. The classification of 3-jets
Same as in Section 5, for $f=(f_{1},f_{2})$ with $\mathrm{ord}(f_{1})=3$, we have $j_{3}(f_{1})\sim x^{3},x^{2}y$ or $x^{3}+xy^{2}$.
@@ 7.1. The classification of 3-jets
II.4 If $a,b,c\neq 0,$ change the notations, $j_{3}(f)\sim(x^{2}y,y^{3}+ax^{3}+bxy^{2}).$ Applying the automorphism $\phi(x)=x,\phi(y)=y+\alpha x$, where $\alpha$ is a nonzero root of $\alpha^{3}-b\alpha^{2}+a=0$ (in this case $\alpha\neq b$ since $a\neq 0$), we have $j_{3}(f)\sim(x^{2}y+\alpha x^{3},y^{3}+(\alpha^{3}+b\alpha^{2}+a)x^{3}+bxy^{2}+2b\alpha x^{2}y)\sim(x^{2}y+\alpha x^{3},y^{3}+bxy^{2}+2b\alpha(x^{2}y+\alpha x^{3}))\sim(x^{2}y+\alpha x^{3},y^{3}+bxy^{2})\sim(x^{3}+x^{2}y,y^{3}+\lambda xy^{2}),\lambda\neq 1$ since $\alpha\neq b$.
III. $j_{3}(f_{1})\sim x^{3}+xy^{2}.$ We have $j_{3}(f)\sim(x^{3}+xy^{2},axy^{2}+bx^{2}y+cy^{3})$. All the discussions is same as that in Section 5, except for the characteristic becomes to $3$, with nothing changes.
@@
$\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},y^{3}+xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{2}y,y^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda xy^{2})(\lambda\neq 1),\\ &(x^{3}+xy^{2},0).\end{split}$ (7.1)
@@ 7.2. The classification of unimodal
$\widetilde{P}_{r,s}$ $(x^{3}+xy^{3}+uxy^{r}+vy^{s},x^{2}y+y^{4}),\ where$ $r\geq 4,\ s\geq 5,$
$u=u_{0}+u_{1}y+\dots,\ v=v_{0}+v_{1}y+\dots,$ $3\mid 2r-2s$
@@
Since most of the discussion is the same as before, we make a table to present the results. See Table LABEL:table_char3-process.
@@
When $j_{3}(f)\sim(x^{3},xy^{2}),(x^{2}y,y^{3}+xy^{2}),(x^{3},x^{2}y)$, we need further discussion.
@@
If $3\mid r$, reset notations, write $f\sim(e(y)x^{3}+y^{r},xy^{2})\sim(x^{3}+e_{1}x^{3}y+y^{r},xy^{2})$, where $e_{1}\in F$. Then $f\sim\widetilde{M}^{3}_{r,e_{1}}$.
@@
If $3\mid r$, similar as above, $f\sim(y^{3}+xy^{2}+\lambda xy^{3}+x^{r},x^{2}y)\sim\widetilde{K}^{3}_{r,\lambda}$ with $\lambda\in F$. ∎
@@
Here we omit the discussion same as Proposition 5.12 to Proposition 5.16 and give the result directly in Table LABEL:eqtable.
@@
$s=4$ $(x^{3}+y^{4},x^{2}y)$ $(4,3;$ $(0,y^{4})$ $(x^{3}+y^{4},x^{2}y+\mu y^{4}),$
$12,11)$ $\mu\in\{0,1\}$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ 0 $(x^{3}+xy^{3},x^{2}y+\lambda y^{4}),$
$r=3$ $\lambda\notin\{0,1\}$ $\lambda\notin\{0,1\}$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ $(0,y^{t}),t\geq 5$ $(x^{3}+xy^{3},x^{2}y+y^{t}),$
$r=3$ $\lambda=0$ $t\geq 5$
$s=5,$ $(x^{3}+xy^{3},x^{2}y+\lambda y^{4})$ $(3,2;9,8)$ $(xy^{r},0),(y^{s},0)$} $(x^{3}+xy^{3}+u(y)xy^{r}$}
$r=3$ $\lambda=1$ $v(y)y^{s},x^{2}y+y^{4})$}
where $u(y),v(y)$ are
units in $F[[x,y]]$,
$r\geq 4,s\geq 5$
@@ 8. Check the modality
Let $T^{1,sec}(f)=R^{m}\bigg/\bigg(\langle f_{1},\dots,f_{m}\rangle\cdot R^{m}+\mathfrak{m}\langle\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}\rangle\bigg)$. Choose $g_{1},\dots,g_{d}$ to be a $F$-basis of $T^{1,sec}(f)$. T.H. Pham, G. Pfister, and G.M. Greuel have shown in [PPG25] that
$F_{\mathbf{t}}(x)=F(x,t_{1},\dots,t_{d})=f+t_{1}g_{1}+\dots+t_{d}g_{d}$
represents a formally semiuniversal deformation of $f$, where $\mathbf{t}=(t_{1},\dots,t_{d})$. If $F(x,t)$ is equivalent to a family of ICIS of at most one parameter for $t\in F^{d}$, then $f$ is unimodal. Here we check $f\sim l_{q,\lambda}\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}),\ q\geq 3,\lambda^{2}\notin\{0,-1\}$ in Table LABEL:table-ord2 as an example.
@@
Note that $T^{1,sec}(f)=(\mathfrak{m}R^{2})/\widetilde{T}_{f}(\mathcal{K}f)$. In the proof of (3.b.ii.2.3.3.1) in Proposition 3.6, we have shown that $(0,xy^{q+1})\in\widetilde{T}_{f}(\mathcal{K}f)$. That means $(0,y^{q+3})\in\widetilde{T}_{f}(\mathcal{K}f)$. We also have $(0,y^{q+2})$ generates $(0,xy^{q})$, $(y^{3},0)$ and $(0,y^{q+1})$ generate $(0,xy^{q-1})$. Then we add $(0,x),(0,xy),\dots,(0,xy^{q-2})$ as generators. These generators form a basis of $T^{1,sec}(f)$.
@@
We consider $F(x,\mathbf{t})=f+t_{1}g_{1}+\dots+t_{2q+1}g_{2q+1}$, where $\mathbf{t}=(t_{1},\dots,t_{2q+1})$. We write $F(x,\mathbf{t})=(G_{1},G_{2})\in R^{2}$.
@@
$G$ $\sim(x^{2}+y^{4},y^{q+2}+\lambda y^{q}+t_{5}y^{4}+\dots+t_{q+3}y^{q+2}+t_{q+5}xy^{2}+\dots+t_{2q+1}xy^{q-2})$ (8.1)
$\sim(x^{2}+y^{4},\sum_{i\geq u}t_{i+1}y^{i}+x\sum_{j\geq v}t_{j+q+3}y^{j}),$
where $q\geq 3,\lambda^{2}\notin\{0,1\},u\geq 4,v\geq 2$. Comparing with (3.1) and Proposition 3.6, it corresponds to the case $\alpha=1,\ s=4,\ u=t\geq 4,\ v=q\geq 2$, where $\alpha,s,t,q$ is in the sense of (3.1).
By Proposition 3.6 (3.a.i), in most of cases, $G$ is at most unimodal. The only unsure case is that there exists $q^{\prime}<t<t^{\prime}$ such that
@@
$\widetilde{l}_{q,t,t^{\prime}}$ $(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+uxy^{t}+xy^{t^{\prime}}),$ $\lambda^{2}=-1,q\geq 3,$ $q\leq 2p+3$
where $u=u_{0}+u_{1}y^{p}+u_{2}y^{2p}+\dots$ $t\geq q+1,t^{\prime}\geq t+1,$
@@
for general $a,b,c,d\in F$, then we can ensure that the singularities given in Table LABEL:table-ord2-mod1 are the only unimodal ICIS. Conversely, if all singularities of the form $\widetilde{l}_{q,t,t^{\prime}}$ are equivalent (or at least can be presented as a one-parameter family), then all the singularities given in Table LABEL:table-ord2 are unimodal. Unfortunately, we cannot judge this equivalence yet. We post it as a conjecture.
@@ II $(x^{3},y^{3}+x^{2}y)$
That means, every singularities in Table LABEL:table-ord3 is unimodal. This is because the bimodular one must have a deformation to $\widetilde{P}_{r,s}$. But $(xy^{k},0),(y^{l},0)\ (k\geq 4,l\geq 5)$ do not exist in a basis of $T^{1,sec}(f)$ at the same time for every $f$. Then such a deformation does not exist.
For ord(f)=3,charF=2ord(f)=3,\ charF=2 case, see Table LABEL:table-ord3-char2-mod1:
@@
$\widetilde{M}_{r}^{2}$ $(x^{3}+y^{r}+ey^{l},x^{2}y+xy^{2}),$ $r\geq 4,\ r\ is\ odd,$ $l\leq 7$
where $e=e_{0}+e_{1}y^{2}+e_{2}y^{4}+\dots$ l>r,lisevenl>r,\ l\ is\ even
@@
For ord(f)=3,char(F)=3ord(f)=3,\ char(F)=3 case, we can check that every singularities in Table LABEL:table-ord3-char3 is unimodal.